# LLM Pruning and Distillation in Practice: The Minitron Approach

### v2

@@ 5 Evaluation
following Touvron et al. [19], we evaluate our compressed models on a series of downstream tasks, including MMLU [20], HumanEval [21] for Python code generation, several question-answering datasets for common-sense reasoning: Arc-C [22], HellaSwag [23], TruthfulQA [24] and WinoGrande [7] and XL-Sum English [25] for summarization. We report the 5-shot performance on MMLU, 5-shot on Winogrande, 25-shot on ARC-Challenge, 10-shot on HellaSwag, 0-shot on 20% of XL-Sum and average pass@1 scores for HumanEval and MBPP. For pass@1 scores we use a temperature of 0.2 and nucleus sampling [26] with top-p === 0.95. For instruction-tuned models, we use MT-Bench [15], Instruction-Following Eval (IFEval) [14], ChatRAG-Bench [16], and Berkeley Function Calling Leaderboard (BFCL) [17].
@@ 5.1 Base Models
[…]
@@ 5.2 Instruct Models
The performance of the instruction-tuned Llama-3.1-Minitron 4B variants is shown in Table 2. We compare the Llama-3.1-Minitron 4B variants to other similarly-sized baselines and notice that our models demonstrate strong instruction-following and roleplay capabilities, only lagging behind Gemma2 in IFEval [14] and MT-Bench [15]. On retrieval based question answering (ChatRAG-Bench [16]) and function-calling (BFCL [17]), Minitron models achieve state-of-the-art performance.
@@ 5.3 Runtime Performance Analysis
We optimized the Llama 3.1 8B and Llama-3.1-Minitron 4B variants with NVIDIA TensorRT-LLM, an open-source toolkit for optimized LLM inference.
Figure 10: TensorRT-LLM FP8 throughput comparison for the Llama-3.1-Minitron-4B models with the Llama 3.1 8B model w.r.t. increasing input and output sequence lengths.
Figure 10 shows the throughput in requests per second for the various models in FP8 precision obtained on a single H100 80 GB GPU. Different use cases are represented by increasing input sequence length/output sequence length (ISL/OSL) combinations, at a batch size of 32 and 64 for the 8B-12B models and the 4B models respectively. The smaller memory footprint of the 4B model allows for larger batches. We notice that Llama-3.1-Minitron-4B-Depth is the fastest, achieving an average throughput improvement of $2.7\times$2.7 × over Llama 3.1 8B; the Llama-3.1-Minitron-4B-Width variant achieves an average throughput improvement of $1.8\times$1.8 × over Llama 3.1 8B. Compared to BF16, we notice that FP8 delivers a performance boost of $1.4\times$1.4 ×.
@@ 6 Insights
In this Section, we summarize some interesting and surprising observations.
6.0.1 General
@@
    In line with the Minitron paper’s observations, we require only 380B tokens to achieve state-of-the-art accuracy post pruning with distillation.
@@
6.0.2 Mistral NeMo 12B to MN-Minitron-8B:
@@
6.0.3 Llama 3.1 8B to Llama-3.1-Minitron 4B:
@@
    Width pruning delivers better accuracy with MMLU at 60.5%, while depth pruning yields 58.7%, for Llama-3.1 compression.
@@
    Reasoning ability is impacted further significantly, with GSM8K accuracy at 41.24% for width and 16.8% for depth.
@@
    Depth pruning boosts throughput, achieving $2.7\times$2.7 × speedup over Llama-3.1 8B, while width pruning provides $1.7\times$1.7 × speedup.

### v3

@@ 5 Evaluation
following Llama [20], we evaluate our compressed base and aligned models on a series of downstream tasks, namely MMLU [21], HumanEval [22] for Python code generation, MBPP [23] and GSM8K [24]. We also evaluate the base models on several question-answering datasets for common-sense reasoning: Arc-C [25], HellaSwag [26], TruthfulQA [27], WinoGrande [10], and XL-Sum English [28] for summarization. The instruction tuned models are further evaluated for question-answering, function calling, instruction following and multiturn conversations on GPQA [29], BFCL [30], IFEval [31] and MT-Bench (GPT4-Turbo) [32], respectively. Note that this MT-Bench is a corrected version of the original MT-Bench [33].
For base models, accuracy is reported with the following evaluations settings: 5-shot on MMLU, 5-shot on Winogrande, 25-shot on ARC-Challenge, 10-shot on HellaSwag, 0-shot on 20% of XL-Sum and average pass@1 scores for HumanEval and MBPP. For pass@1 scores we use a temperature of 0.2 and nucleus sampling [34] with top-p $=$ 0.95. For aligned models we use 0 shot and greedy sampling if applicable.
@@ 5.1 Base Models
Base model evaluation results are shown in Table 1. Compared to similarly-sized models, MN-Minitron-8B demonstrates superior accuracy across the board, outperforming the recent Llama 3.1 8B model using 40$\times$ fewer training tokens (380B vs. 15T). Similarly, the Llama 3.1-Minitron-4B models perform favorably compared to the teacher Llama 3.1 8B model using 150×150$\times$150 × fewer training tokens (94B vs. 15T); our pruned Llama models also outperform the original Minitron 4B model [2]. We note from Table 1 that the width-pruned Llama variant outperforms the depth-pruned one. These results clearly demonstrate the advantages of our methodology: state-of-the-art accuracy coupled with an order of magnitude improvement in training efficiency.
@@ 5.2 Instruct Models
The accuracy of the instruction-tuned model variants are shown in Table 2. Our aligned models outperform similarly sized variants on most evaluated benchmarks with the exception of HumanEval [35] and MBPP [23]. Additionally, Llama 3.1-Minitron-4B lags behind Gemma2 on MT-Bench [33]. Nevertheless, our aligned models are consistently better on MMLU [21], GSM8K [24], GPQA [29], IFEval [31] and BFCLv2 [30]. This demonstrates the strong capabilities of our model.
@@ 5.3 Runtime Performance Analysis
To evaluate runtime performance, we optimize the Llama 3.1 8B and Llama 3.1-Minitron-4B variants with NVIDIA TensorRT-LLM, an open-source toolkit for optimized LLM inference.
Figure 10: TensorRT-LLM FP8 throughput comparison for the Llama 3.1-Minitron-4B models with the Llama 3.1 8B model w.r.t. increasing input and output sequence lengths.
Figure 10 shows the throughput in requests per second for the various models in FP8 precision obtained on a single H100 80 GB GPU. Different use cases are represented by increasing input sequence length/output sequence length (ISL/OSL) combinations, at a batch size of 32 and 64 for the 8B-12B models and the 4B models respectively. The smaller memory footprint of the 4B model allows for larger batches. We notice that Llama 3.1-Minitron-4B (Depth) is fastest, achieving an average throughput improvement of 2.7×2.7$\times$2.7 × over Llama 3.1 8B; the width-pruned variant achieves an average throughput improvement of 1.8×1.8$\times$1.8 × over Llama 3.1 8B. Compared to BF16, we notice that FP8 delivers a performance boost of 1.4×1.4$\times$1.4 ×.
@@ 6 Insights
In this section, we summarize some interesting and surprising observations based on our evaluation.
General
@@
    In line with the Minitron paper’s observations, we require a order of magnitude fewer tokens (380B vs 15T) to achieve state-of-the-art accuracy post pruning with distillation.
@@
Mistral NeMo 12B to MN-Minitron-8B
@@
Llama 3.1 8B to Llama 3.1-Minitron-4B
@@
    Width pruning delivers better accuracy with MMLU at 60.5%, while depth pruning yields 58.7%, for Llama 3.1 compression.
@@
    Reasoning ability for base variants appears to be impacted significantly for the depth pruned version, with GSM8K accuracy at 16.8% compared to 41.24% for the width pruned version. However, the gap reduces with instruct tuning.
@@
    Depth pruning boosts throughput, achieving 2.7×2.7$\times$2.7 × speedup over Llama-3.1 8B, while width pruning provides 1.7×1.7$\times$1.7 × speedup.